# Lessons Learned So Far From Verifying the Rust Standard Library (work-in-progress)

### v1

@@ 2.1. Verification progress
Challenge (#) Progress Tool used Reward (USD)
@@
SIMD intrinsics (15) Complete Randomized testing* 20,000
@@
    The results of randomized test cases involving models of the SIMD intrinsics and actual intrinsics were compared
@@ 6 Some(x * 2)
A second backend for Kani is the goto-transcoder, which generates inputs suitable for the ESBMC tool (Cordeiro et al., 2009) (rather than CBMC, as usually targetted by Kani). Using the goto-transcoder adds support for $k$-induction, SMT solvers, and concurrency models. In principle, this could result in more powerful verification (more cases verified) using less CPU time and memory, but, for now, no challenges have been solved using goto-transcoder rather than Kani.
@@ 2.3. Suggested tools for remaining challenges
    Does it require reasoning in a concurrent context? Accepted tools with well-documented support for this: VeriFast, goto-transcoder.
@@
atomic types & intrinsics (7) $\times$ $\times$ VF, goto-t
@@
Arc (27) $\times$ $\times$ VF, goto-t
  • ‘VF’ refers to VeriFast, ‘goto-t’ refers to goto-transcoder.
@@ 4.3. Verification of concurrent code
Although safe Rust holds out the promise of “fearless concurrency” for its users, the promise does not extend to unsafe Rust. It is possible to write unsafe Rust that contains race conditions, which are immediate undefined behaviour in Rust. Thus, verifying the standard library must eventually include some provision for reasoning about concurrency, where appropriate. In fact, of the existing challenges, two explicitly require verifying an absence of data races (Challenges 7 and 27). Having said that, if unsafety is encapsulated sufficiently that the unique ownership property holds, then it should be sound to verify relevant methods as if they were in a sequential program.

### v2

@@ 2.1. Verification progress
Challenge (#) Progress* Tool used Reward (USD)
@@
SIMD intrinsics (15) Complete Randomized testing** 20,000
@@
    Completed by: (1) Alex Le Blanc, Patrick Lam (University of Waterloo) – (3) Surya Togaru, Yifei Wang, Szu-Yu Lee, Mayuresh Joshi (CMU) – (5) Bart Jacobs (KU Leuven) – (6) Quinyuan Wu, Daniel Tu, Dhvani Kapadia, Jiun Chi Yang (CMU) – (9) Samuel Thomas, Cole Vick (Unknown) – (11) Rajath M Kotyal, Yen-Yun Wu, Lanfei Ma, Junfeng Jin (CMU) – (14) Shoyu Vanilla (Unknown) – (15) Aniket Mishra, Karthikeyan Bhargavan, Maxime Buyse (Cryspen) – (19) Bart Jacobs (KU Leuven).
    The results of randomized test cases involving models of the SIMD intrinsics and actual intrinsics were compared.
@@ 6 Some(x * 2)
A second backend for Kani is the goto-transcoder, which generates inputs suitable for the ESBMC tool (Cordeiro et al., 2009) (rather than CBMC, as usually targetted by Kani). Using the goto-transcoder adds support for $k$-induction and SMT solvers. In principle, this could result in more powerful verification (more cases verified) using less CPU time and memory, but, for now, no challenges have been solved using goto-transcoder rather than Kani.
@@ 2.3. Suggested tools for remaining challenges
    Does it require reasoning in a concurrent context? Accepted tools with well-documented support for this: VeriFast.
@@
atomic types & intrinsics (7) $\times$ $\times$ VF
@@
Arc (27) $\times$ $\times$ VF
  • ‘VF’ refers to VeriFast.
@@ 4.3. Verification of concurrent code
Although safe Rust holds out the promise of “fearless concurrency” for its users, the promise does not extend to unsafe Rust. It is possible to write unsafe Rust that contains data race conditions, which are immediate undefined behaviour in Rust. Thus, verifying the standard library must eventually include some provision for reasoning about concurrency, where appropriate. In fact, of the existing challenges, two explicitly require verifying an absence of data races (Challenges 7 and 27). Having said that, if unsafety is encapsulated sufficiently that the unique ownership property holds, then it should be sound to verify relevant methods as if they were in a sequential program.